What type is this following
8y=7x
-2x=6y
is it independent
dependent
inconsistent
or none of these?

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4169)

inconsistent,because there is no solution to this problem

Answer by Alan3354(31538)

What type is this following
8y=7x
-2x=6y
is it independent
dependent
----------
7x-8y = 0
-2x-6y = 0 --> x + 3y = 0 times 7 --> 7x + 21y = 0
-----
7x + 21y = 0
7x - 8y = 0
-------------- Subtract
29y = 0
y = 0
-------
x = 0
(0,0) is the solution --> independent